An architect designs a rectangular flower garden, the width is exactly 2/3's of the length. If 260 feet of picket fencing r to b used to enclose the garden, find the dimension of the garden...HELP

Answer by shree840(260)

let length be x
width = (2/3)x
Perimeter = x+x+(2/3)x(2/3)x=260 ft
=2x+(4/3)x=260
6x+4x=260*3=780 (Multiplying both sides by 3 to get rid of denominator
or 10x =780
x=78
width =2/3* 78=52
check 78+78+52+52= 260